# backgroundcolor=, commentstyle=, breakatwhitespace=false, keywordstyle=, language=Java, stringstyle=, basicstyle=, showstringspaces=false

### v1

@@ 5.2 RQ2: Does the inclusion of source code help to enhance the detection of technical debt?
Motivation: As mentioned earlier, existing work on TD detection has primarily focused on comments and other textual artifacts such as commits and issues [42, 16, 37, 36, 55] while overlooking the content of source code. However, certain features of source code can indicate the presence of some types of technical debt, such as design or implementation. For instance, design technical debt may be introduced in anti-pattern source code where developers have neglected specific design principles [76]. Recent studies have established a link between code smells and technical debt [67]. This highlights the potential of source code to provide meaningful features for the detection of technical debt. Furthermore, during the annotation process, it was observed that annotators were more proficient in identifying technical debt when utilizing source code information as a reference, rather than relying exclusively on comments. Thus, our objective is to explore the effectiveness of leveraging source code for the SATD detection task.
TABLE V: Performance (F1-score) comparison of various models on SATD detection using only comments versus incorporating additional source code. The subscripts accompanying the numerical results indicate the number of context lines that produced the best outcomes for each model, with ff representing the use of the full function.
Model Comment Comment + Code only StrConcat CodeAtt Ensemble RoBERTa 68.28 $69.11_{2}$ ( $\uparrow$1.22%) $69.45_{2}$ ( $\uparrow$1.71%) 69.92 CodeBERT 72.75 $76.30_{\text{ff}}$ ( $\uparrow$4.88%) $74.80_{\text{ff}}$ ( $\uparrow$2.82%) 76.55 UniXCoder 70.62 $71.54_{20}$ ( $\uparrow$1.30%) $71.46_{\text{ff}}$ ( $\uparrow$1.19%) 72.22 GraphCodeBERT 71.39 $74.86_{2}$ ( $\uparrow$4.86%) $72.75_{\text{ff}}$ ( $\uparrow$1.91%) 75.25
@@
  • Model versatility: Our goal is to explore the impact of incorporating source code in various models to demonstrate the effectiveness of this approach. As such, we apply the proposed method to four PLMs: RoBERTa, CodeBERT, UniXCoder and GraphCodeBERT. BERT has been excluded because it showed performance similar to or lower than its variant, RoBERTa as can be seen in Table IV.
@@
      String Concatenation (StrConcat): Two sequences of tokens are concatenated then passed into pre-trained models. This approach is typically employed in prior studies for text classification tasks [17, 39, 31].
@@
      $\mathcal{A}=softmax(G\cdot H^{T})$
      $classification\_emb=\mathcal{A}\cdot H^{T}$
  • Code context length: Based on our observations during the labeling process, annotators did not need to review the entire code to determine the type of technical debt (TD) associated with a comment; they only needed to scan the nearby code. With this in mind, we investigated the impact of varying the length of the code context. Specifically, we assessed the effect of using the surrounding code by including 2, 10, and 20 lines, as well as the entire function. Figure 2 illustrates an example of utilizing a code context that includes 2 lines of code.
Result: Table V presents the performance of the four PLMs in detecting SATD comments, comparing the outcomes of using only comments versus combining comments with code context. The analysis of the results is based on the three previously mentioned aspects.
@@
  • Integration techniques: Overall, both proposed methods improve performance across all models, with StrConcat demonstrating greater effectiveness than CodeAtt except in the case of RoBERTa. Specifically, StrConcat enhances the performance of CodeBERT and GraphCodeBERT by over 4.88% and 4.86%, respectively. In contrast, the improvements achieved with CodeAtt are more moderate, ranging between 1.19% and 2.82%. The superior results of StrConcat highlight the ability of Transformer-based models to effectively process multi-modal inputs through their self-attention mechanism, a capability that has been demonstrated in various downstream tasks [73, 69, 83].
@@
  • Code context length: Table V shows that the F1-Score of four models is significantly improved when using either two lines of code context or the entire function code. The tendency to favor two surrounding code lines during prediction aligns with human intuition, which relies on local context for annotation. Moreover, incorporating the entire function code highlights the capability of these models to leverage global context. This context could enhance the performance by enabling models to identify relevant code snippets across the function, offering a more nuanced understanding of the function’s structure and semantic, thereby potentially improving the reliability of the models. While employing code context generally demonstrates improvements over using comments alone, varying the context length might impact model performance differently. Consequently, we employ an ensemble approach to combine model predictions across different code context lengths. Specifically, for each model a majority voting mechanism was applied to produce the final prediction. Each model was configured with varying code context lengths, including 2, 10, 20 lines and the entire function code, considering CodeAtt as the input concatenation approach for RoBERTa and StrConcat for the other models. As shown in the last column of Table V, this ensemble approach achieves the highest performance across all four models, underscoring the advantage of leveraging multiple code context lengths for identifying SATD comments.
@@
Answer to RQ2 Incorporating comments with source code information results in performance improvements across various models compared to using comments alone, highlighting the robustness, versatility, and adaptability of this approach in detecting SATD. The proposed methods, StrConcat and CodeAtt, effectively utilize the source code context and enhance the performance across all evaluated models. This paves the way for future research with the ultimate aim of further improving the prediction. When comment and code context are combined as input, an optimal performance is achieved with 2 surrounding code lines or by including the entire code function. Combining various scopes demonstrates the effective contribution of each scope, highlighting the potential of multi-code scope strategies in improving SATD detection.
@@
5.3 RQ3: What is the accuracy of different pre-trained models when detecting TD solely from source code?
@@
Motivation: As mentioned earlier, technical debt is frequently identified through textual content, such as comments or issue reports. However, when such debt is not explicitly documented, existing tools are unable to detect it, despite its presence in the code. Moreover, many comments become outdated or inconsistent with the actual code, as developers often fail to update comments after modifying the code. This discrepancy between comments and code introduces a significant blind spot for tools that rely solely on textual indicators, limiting their ability to accurately detect technical debt. Therefore, there is a pressing need for more advanced approaches that surpass textual cues to effectively identify and manage technical debt within code bases.
@@
TABLE VI: Performance of different PLMs on TD detection using Tesorocode.
Model Model size EM F1 Encoder-based PLMs CodeBERT [22] 125M 38.28 43.47 UniXCoder [23] 125M 38.12 42.58 GraphCodeBERT [24] 125M 39.38 44.21 RoBERTa [39] 125M 35.37 38.22 ALBERT [31] 11.8M 39.32 41.99 Encoder-Decoder-based PLMs PLBART [2] 140M 36.85 39.90 Codet5 [71] 220M 32.66 35.41 CodeT5+ [70] 220M 37.91 41.96 Decoder-based PLMs (LLMs) TinyLlama [79] 1.03B 37.05 40.05 DeepSeek-Coder [82] 1.28B 42.52 46.19 OpenCodeInterpreter [25] 1.35B 38.16 41.76 phi-2 [54] 2.78B 37.92 41.57 starcoder2 [40] 3.03B 35.37 41.77 CodeLlama [54] 6.74B 34.14 38.16 Magicoder [75] 6.74B 39.14 42.49
@@
We conducted experiments on 16 models, including 5 Encoder-based PLMs, 3 Encoder-Decoder-based PLMs, and 8 Decoder-based PLMs. Since only code is used as input, and Sections 5.1 and 5.2 demonstrated the superior performance of code-based PLMs, we experiment with models primarily pre-trained on coding corpora. The language model head layer is replaced with a linear classification head during fine-tuning on the downstream task. All models are fine-tuned for 10 epochs using a batch size of 32 and a learning rate of $51e-$ 5. For models with more than 6 billion parameters, we utilize LoRA [27] with a learning rate of $31e-$ 3 for fine-tuning due to resource constraints. For decoder-base PLMs, we apply a template for input presented in the online appendix, and use the embedding of final tokens as the representation fed into the classification head. We randomly split Tesoro into 10 folds for cross-validation and report the average Exact Match (EM) and F1-score across these folds for each model.
@@
Result: Table VI shows the experimental results, with bold text indicating the highest score, while underlined scores representing the runner-up. DeepSeek-Coder achieves the best performance with an F1-score of 46.19% marking an improvement of 4.98% over the second highest one, i.e., GraphCodeBERT getting 44.21%. Though previous studies indicated a limited adaptability of LLMs to classification tasks [80, 63], these results highlight the potential of such models. However, models containing the Decoder module generally exhibit lower performance compared to Encoder-based models. Figure 7 supports this observation, as the three models following DeepSeek-Coder are all Encoder-based. Several factors can account for this observation. Firstly, Encoder-Decoder and Decoder-based models are pretrained on generation tasks, which may result in suboptimal performance on classification tasks due to the lack of task-specific optimization. Secondly, some studies [5, 34] showed that Decoder-based models are less effective for text representation or embeddings due to their causal attention mechanism, which limits the model’s ability to learn robust representations. Hence, the embedding information before the classification head is not sufficiently rich, leading to a suboptimal performance. Though Encoder-based models experience a slight performance drop compared to that of DeepSeek-Coder, they achieve this with significantly fewer parameters–around 90% less–offering a more practical approach to TD detection using source code.
@@
Answer to RQ3 DeepSeek-Coder achieves the highest accuracy on Tesorocode. In contrast, Encoder-based models exhibit a slight performance drop but with a substantial reduction in parameters, making them more practical in real-world scenarios. Code-based PLMs show superiority in the detection of TD from source code. However, their performance remains below 50% in both EM and F1-score. This indeed highlights the need for more advanced methods and further research.
6 Discussion

### v2

@@ 5.1. RQ1: Does the inclusion of source code help to enhance the detection of technical debt?
Motivation: As mentioned earlier, most existing work on TD detection has primarily focused on comments and other textual artifacts such as commits and issues (Maldonado and Shihab, 2015; da Silva Maldonado et al., 2017; Li et al., 2023, 2020; Sala et al., 2021) while overlooking the content of source code. However, certain features of source code can indicate the presence of some types of technical debt, such as design or implementation. For instance, design technical debt may be introduced in anti-pattern source code where developers have neglected specific design principles (Xiao et al., 2022). This highlights the potential of source code to provide meaningful features for the detection of technical debt. Furthermore, during the annotation process, it was observed that annotators were more proficient in identifying technical debt when utilizing source code information as a reference, rather than relying exclusively on comments. A few recent studies have tried to leverage source code alongside comments to improve SATD detection, demonstrating that code context can enhance model performance (Russo et al., 2022; Zampetti et al., 2020; Russo et al., 2025). However, these approaches often lack a comprehensive analysis into critical factors such as the effect of different code-comment integration techniques or the optimal scope of code context–instead frequently relying on simplistic strategies of using full code. Therefore, this RQ aims to evaluate the effectiveness of incorporating source code into SATD detection and to systematically investigate various factors to identify the most effective strategies for utilizing code context.
@@
  • Model versatility: Our goal is to explore the impact of incorporating source code in various models to demonstrate the effectiveness of this approach. As such, we apply the proposed method to four PLMs: RoBERTa, CodeBERT, UniXCoder and GraphCodeBERT. BERT has been excluded because it showed performance similar to or lower than its variant, RoBERTa as can be seen in Table 6.
@@
      String Concatenation (StrConcat): Two sequences of tokens are concatenated then passed into pre-trained models. This approach is typically employed in prior studies for text classification tasks (Devlin et al., 2018; Liu et al., 2019; Lan et al., 2019).
@@
      $A=softmax(G\cdot H^{T})$
      $classification\_emb=A\cdot H^{T}$
  • Code context scope: Based on our observations during the labeling process, annotators did not need to review the entire code to determine the type of technical debt (TD) associated with a comment; they only needed to scan the nearby code. With this in mind, we investigated the impact of varying the length of the code context. Specifically, we assessed the effect of using the surrounding code by including 2, 10, and 20 lines, as well as the entire function. Figure 2 illustrates an example of utilizing a code context that includes 2 lines of code.
Table 4. Performance (F1-score) comparison of various models on SATD detection using only comments versus incorporating additional source code. The subscripts accompanying the numerical results indicate the number of context lines that produced the best outcomes for each model, with $f$ representing the use of the full function.
Model Comment Comment + Code
only StrConcat CodeAtt Ensemble
RoBERTa 68.28 $69.11_{2}$ ( $\uparrow$1.22%) $69.45_{2}$ ( $\uparrow$1.71%) 69.92
CodeBERT 72.75 $76.30_{\text{ff}}$ ( $\uparrow$4.88%) $74.80_{\text{ff}}$ ( $\uparrow$2.82%) 76.55
UniXCoder 70.62 $71.54_{20}$ ( $\uparrow$1.30%) $71.46_{\text{ff}}$ ( $\uparrow$1.19%) 72.22
GraphCodeBERT 71.39 $74.86_{2}$ ( $\uparrow$4.86%) $72.75_{\text{ff}}$ ( $\uparrow$1.91%) 75.25
Result: Table 4 presents the performance of the four PLMs in detecting SATD comments, comparing the outcomes of using only comments versus combining comments with code context. The analysis of the results is based on the three previously mentioned aspects.
@@
  • Integration techniques: Overall, both proposed methods improve performance across all models, with StrConcat demonstrating greater effectiveness than CodeAtt except in the case of RoBERTa. Specifically, StrConcat enhances the performance of CodeBERT and GraphCodeBERT by over 4.88% and 4.86%, respectively. In contrast, the improvements achieved with CodeAtt are more moderate, ranging between 1.19% and 2.82%. The superior results of StrConcat highlight the ability of Transformer-based models to effectively process multi-modal inputs through their self-attention mechanism, a capability that has been demonstrated in various downstream tasks (Wei et al., 2020; Wang et al., 2022; Zhou et al., 2023).
@@
  • Code context scope: Table 4 shows that the F1-Score of four models is significantly improved when using either two lines of code context or the entire function code. The tendency to favor two surrounding code lines during prediction aligns with human intuition, which relies on local context for annotation. Moreover, incorporating the entire function code highlights the capability of these models to leverage global context. This context could enhance the performance by enabling models to identify relevant code snippets across the function, offering a more nuanced understanding of the function’s structure and semantic, thereby potentially improving the reliability of the models. While employing code context generally demonstrates improvements over using comments alone, varying the context length might impact model performance differently. Consequently, we employ an ensemble approach to combine model predictions across different code context lengths. Specifically, for each model a majority voting mechanism was applied to produce the final prediction. Each model was configured with varying code context lengths, including 2, 10, 20 lines and the entire function code, considering CodeAtt as the input concatenation approach for RoBERTa and StrConcat for the other models. As shown in the last column of Table 4, this ensemble approach achieves the highest performance across all four models, underscoring the advantage of leveraging multiple code context lengths for identifying SATD comments.
@@
Answer to RQ1 Incorporating comments with source code information results in performance improvements across various models compared to using comments alone, highlighting the robustness, versatility, and adaptability of this approach in detecting SATD. The proposed methods, StrConcat and CodeAtt, effectively utilize the source code context and enhance the performance across all evaluated models. This paves the way for future research with the ultimate aim of further improving the prediction. When comment and code context are combined as input, an optimal performance is achieved with 2 surrounding code lines or by including the entire code function. Combining various scopes demonstrates the effective contribution of each scope, highlighting the potential of multi-code scope strategies in improving SATD detection.
@@
5.2. RQ2: What is the accuracy of different pre-trained models when detecting TD solely from source code?
@@
Motivation: As mentioned earlier, technical debt is frequently identified through textual content, such as comments or issue reports. However, when such debt is not explicitly documented, existing tools are unable to detect it, despite its presence in the code. Moreover, many comments become outdated or inconsistent with the actual code, as developers often fail to update comments after modifying the code. This discrepancy between comments and code introduces a significant blind spot for tools that rely solely on textual indicators, limiting their ability to accurately detect technical debt. Besides, some prior studies have explored technical debt detection directly within source code; however, they typically focused on limited classification schemes–such as distinguishing only between high and not high TD (Tsoukalas et al., 2024, 2022), or on identifying the mere presence or absence of code smells (Yadav et al., 2024), without capturing the broader diversity of technical debt types. Therefore, there is a pressing need for more advanced approaches that surpass textual cues to effectively identify and manage technical debt within code bases. In response, we propose investigation on detecting multi-type TD in source code, extending beyond conventional binary classification frameworks.
@@
Table 5. Performance of different PLMs on TD detection using Tesorocode.
Model Model size EM F1
Encoder-based PLMs
CodeBERT (Feng et al., 2020) 125M 38.28 43.47
UniXCoder (Guo et al., 2022) 125M 38.12 42.58
GraphCodeBERT (Guo et al., 2020) 125M 39.38 44.21
RoBERTa (Liu et al., 2019) 125M 35.37 38.22
ALBERT (Lan et al., 2019) 11.8M 39.32 41.99
Encoder-Decoder-based PLMs
PLBART (Ahmad et al., 2021) 140M 36.85 39.90
Codet5 (Wang et al., 2021) 220M 32.66 35.41
CodeT5+ (Wang et al., 2023) 220M 37.91 41.96
Decoder-based PLMs (LLMs)
TinyLlama (Zhang et al., 2024b) 1.03B 37.05 40.05
DeepSeek-Coder (Zheng et al., 2024) 1.28B 42.52 46.19
OpenCodeInterpreter (Guo et al., 2024) 1.35B 38.16 41.76
phi-2 (Roziere et al., 2023) 2.78B 37.92 41.57
starcoder2 (Lozhkov et al., 2024) 3.03B 35.37 41.77
CodeLlama (Roziere et al., 2023) 6.74B 34.14 38.16
Magicoder (Wei et al., 2023) 6.74B 39.14 42.49
@@
We conducted experiments on 16 models, including 5 Encoder-based PLMs, 3 Encoder-Decoder-based PLMs, and 8 Decoder-based PLMs. Since only code is used as input, and Sections 5.1 and 5.2 demonstrated the superior performance of code-based PLMs, we experiment with models primarily pre-trained on coding corpora. The language model head layer is replaced with a linear classification head during fine-tuning on the downstream task. All models are fine-tuned for 10 epochs using a batch size of 32 and a learning rate of 1e51e-5. For models with more than 6 billion parameters, we utilize LoRA (Hu et al., 2021) with a learning rate of 1e31e-3 for fine-tuning due to resource constraints. For decoder-base PLMs, we apply a template for input presented in the online appendix, and use the embedding of final tokens as the representation fed into the classification head. We randomly split Tesoro into 10 folds for cross-validation and report the average Exact Match (EM) and F1-score across these folds for each model.
@@
Result: Table 5 shows the experimental results, with bold text indicating the highest score, while underlined scores representing the runner-up. DeepSeek-Coder achieves the best performance with an F1-score of 46.19% marking an improvement of 4.98% over the second highest one, i.e., GraphCodeBERT getting 44.21%. Though previous studies indicated a limited adaptability of LLMs to classification tasks (Zhang et al., 2024a; Sun et al., 2023), these results highlight the potential of such models. However, models containing the Decoder module generally exhibit lower performance compared to Encoder-based models. Figure 6 supports this observation, as the three models following DeepSeek-Coder are all Encoder-based. Several factors can account for this observation. Firstly, Encoder-Decoder and Decoder-based models are pretrained on generation tasks, which may result in suboptimal performance on classification tasks due to the lack of task-specific optimization. Secondly, some studies (BehnamGhader et al., 2024; Lee et al., 2024) showed that Decoder-based models are less effective for text representation or embeddings due to their causal attention mechanism, which limits the model’s ability to learn robust representations. Hence, the embedding information before the classification head is not sufficiently rich, leading to a suboptimal performance. Though Encoder-based models experience a slight performance drop compared to that of DeepSeek-Coder, they achieve this with significantly fewer parameters–around 90% less–offering a more practical approach to TD detection using source code.
@@
Answer to RQ2 DeepSeek-Coder achieves the highest accuracy on Tesorocode. In contrast, Encoder-based models exhibit a slight performance drop but with a substantial reduction in parameters, making them more practical in real-world scenarios. Code-based PLMs show superiority in the detection of TD from source code. However, their performance remains below 50% in both EM and F1-score. This indeed highlights the need for more advanced methods and further research.
5.3. RQ3: How do the manually classified comments contribute to the detection of SATD?
Table 6. The performance (F1-score) of five PLMs across three tasks when trained on the M-62K (Maldonado-62K) dataset and further enhanced with the additional Tesoro dataset.
Model Identification Classification Detection
M-62K +Tesoro ($\Delta$) M-62K +Tesoro ($\Delta$) M-62K +Tesoro ($\Delta$)
BERT 87.96 88.99 ( $\uparrow$1.17%) 51.42 55.10 ( $\uparrow$7.16%) 45.99 49.64 ( $\uparrow$7.94%)
RoBERTa 89.06 89.96 ( $\uparrow$1.01%) 53.91 55.32 ( $\uparrow$2.62%) 46.13 52.86 ( $\uparrow$14.59%)
UniXCoder 88.38 88.42 ($\uparrow$0.05%) 54.82 54.99 ( $\uparrow$0.31%) 50.94 52.11 ( $\uparrow$2.30%)
CodeBERT 88.74 90.06 ( $\uparrow$1.49%) 57.50 63.70 ( $\uparrow$10.78%) 53.79 55.60 ( $\uparrow$3.36%)
GraphCodeBERT 89.94 90.12 ( $\uparrow$0.20%) 58.00 60.44 ( $\uparrow$4.21%) 49.12 56.87 ( $\uparrow$15.78%)
Motivation: Our primary focus in this RQ is to assess whether the newly identified SATD comments from the Stack corpus contribute positively to the performance of SATD comment detection. Our objective is to enhance the state-of-the-art benchmark dataset introduced by da Silva Maldonado et al. (da Silva Maldonado et al., 2017), henceforth referred to as Maldonado-62K. This dataset comprises 62,566 comments, of which 4,071 (approximately 6.5%) encompass one of five categories of technical debt. While the dataset’s volume is relatively limited, making it challenging to apply deep learning models, it is crucial to increase the diversity of data samples within each category to enhance the training performance of SATD detection models.
Approach: For this RQ, our objective is to examine whether manually classified comments from the Tesorocomment dataset improve the detection of SATD comments. Following previous studies (da Silva Maldonado et al., 2017; Ren et al., 2019; Sala et al., 2021; Sheikhaei et al., 2024), we employed a cross-project experimental approach. The benchmark dataset Maldonado-62K comprises 10 projects, divided into 10 validation folds. We used one fold (one project) for testing and the remaining nine folds (nine other projects) for training. To prevent data leakage, we removed duplicate entries from the entire dataset before splitting it, resulting in 38,269 samples. We then analyzed the impact of incorporating Tesoro during training, denoted as ++Tesoro, compared to using only the state-of-the-art dataset, Maldonado-62K, across three scenarios: SATD-Identification, SATD-Classification and SATD-Detection.
  1. (1)
  2. (2)
  3. (3)
For each scenario, we used the same test set while training the model separately with the Maldonado-62K and with ++Tesoro datasets. Research in the field of SATD detection has highlighted the promising results of PLMs (Di Salle et al., 2022; Sharma et al., 2022; Sheikhaei et al., 2024), thus we also employed these models to detect SATD comments. Specifically, we experimented with BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), CodeBERT (Feng et al., 2020), UniXCoder (Guo et al., 2022) and GraphCodeBERT (Guo et al., 2020). PLMs are employed as comment encoders, followed by a fully connected network dedicated to downstream tasks such as identification, classification, and detection, as previously described. All models are fine-tuned for 10 epochs using a batch size of 32 and a learning rate of 1e51e-5. Given the significant class imbalance in the dataset, the F1 score to was used to evaluate performance.
Result: Table 6 depicts the performance of five different PLMs across three aforementioned scenarios $\mathbb{S}_{1}$, $\mathbb{S}_{2}$ and $\mathbb{S}_{3}$. We see that incorporating Tesoro during the training phase consistently boosts the performance of all PLMs across all tasks. Specifically, the identification of SATD comments from the entire set of comments demonstrates an improvement ranging from 0.05% to 1.49% across all models when the training phase is supplemented with the Tesoro dataset, as opposed to relying solely on the Maldonado-62K dataset. In this scenario, the dataset demonstrates a considerable imbalance, with non-SATD data making up more than 90% of the total. The improvement highlights the effectiveness of the Tesoro comment in mitigating the data imbalance issue, and this is significant because real-world scenarios will have a similar imbalance. In the context of SATD detection, BERT and RoBERTa exhibited lower performance compared to the other three code-based PLMs. However, training with ++Tesoro considerably improves the performance of these two models, leading to comparable results across all the models. For instance, applying ++Tesoro during training improved BERT’s performance by 7.94% and RoBERTa’s by 14.59% in detecting various types of SATD comments. A similar trend is observed in other scenarios, where all models show enhanced performance when utilizing ++Tesoro. Specifically, the performance in classifying the five SATD types increases by approximately 0.31% to 10.78% for all PLM models. Additionally, the results indicate that CodeBERT and its variant, GraphCodeBERT, achieve the highest performance across all scenarios, highlighting the advantage of code-based PLMs for the detection of SATD comments.
Figure 7. An in-depth analysis of CodeBERT and RoBERTa performance across three scenarios for 10 projects.
Finally, we conducted an in-depth analysis of the results from 10 large-scale open source projects across all three scenarios. In this setup, each project serves as a test fold, while the PLMs are trained on the remaining 9 projects. For illustrative purposes, we present the findings for CodeBERT and RoBERTa, as depicted in Figure 7. Overall, it is evident that training the models with ++Tesoro results in improved performance across all projects. In particular, in the $\mathbb{S}_{1}$ task, the performance improved across all tested projects for both models, highlighting the enhanced training set’s effectiveness in addressing the imbalance issue in the original dataset. Furthermore, 7 out of the 10 folds show an increase in F1-score when training CodeBERT on classification and detection tasks with ++Tesoro, while RoBERTa exhibits significant improvement in 9 out of 10 test sets for the SATD detection task.
Answer to RQ3 There is an improvement in the prediction performance when Tesoro is incorporated into the training, validating the efficacy of our data pipeline in selecting informative samples, and proving a robust annotation process. CodeBERT and GraphCodeBERT consistently achieve superior performance and notable improvements across all three tasks, highlighting the advantages of employing code-based PLMs for SATD comment detection.
6. Discussion